OpenStudy (anonymous):

The illumination of the light varies inversely as the as the square of its distance from the light. If the light has an illumination of 100 candle when the distance from the light is 2 feet, what is the lights illumination when the distance from the light is 5 feet?

OpenStudy (anonymous):

Hint: 2^2=4 and 5^2=25

OpenStudy (anonymous):

How should I work this problem out

OpenStudy (anonymous):

I need the answere please

OpenStudy (anonymous):

Whats the answere please I need to turn this in in 20min

OpenStudy (anonymous):

Your constant is K/4=100. So what does K/25=? Find K and you have your answer.

OpenStudy (anonymous):

Thanks